Suppose you are located 180 m from a radio transmitter.What if this station is 98.1 FM? (The FM band frequencies are in megahertz.)


1
Expert's answer
2021-02-22T10:19:24-0500

Radio waves are the electromagnetic waves, so the velocity of the radio waves is "3\times 10^{8}m/s" .

The frequency of "98.1FM"  band (in megahertz) is "98.1 \times 10^6 Hz" .

Now, Substitute  "98.1\times10^6 Hz" for "f"  and "3.0 \times10^8 m/s"  for "v" in the equation "\lambda = \dfrac{v}{f}"

"\lambda= \dfrac{3.0\times10^8m/s}{98.1\times10^6Hz}=3.058m"

Now, Substitute 3.058m for "\lambda" in the equation, "n\lambda=180" m

"\because" "n\lambda = 180"

So, "n= \dfrac{180}{\lambda}"


"= \dfrac{180}{3.058}\ =58.56"

Round off to the three significant digits,

"n=58.9"

Hence, the number of wavelengths present between radio station and you are 58.9 or 60(approx.)